OpenStudy (anonymous):

Simplify 5 log2 k − 8 log2 m + 10 log2 n. 7 log2 (k − m + n) 7 log2 kn over m log2 50 kn over 8 m log2 k to the fifth power n to the tenth power over m to the eighth power

OpenStudy (anonymous):

@uri @LeAnn01Xxx @e.mccormick @Lady.Liv1776 @TylerMJ325 @Terrorclaw @Phebe @keshaun455 @JoshKoikkara

OpenStudy (anonymous):

Hmm... 5*log2 k - 8*log2 m + 10*log2 n log2 (k^5*n^10)/m^8 , I'm not sure...do u get the same?

OpenStudy (anonymous):

I don't want the answer I want to learn how to do it. I'm really bad at Maths ...

Nnesha (nnesha):

quotient rule\[\large\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\] remember!

OpenStudy (anonymous):

@Marcos_RG4L Hmm, I forgot Latex wait.... http://openstudy.com/study#/updates/4ffb5ce1e4b00c7a70c465e9

OpenStudy (anonymous):

@Nnesha Hmm...Great !!

Nnesha (nnesha):

first apply power property \[\large\rm log_b x^y = y \log_b x\]

OpenStudy (anonymous):

5 log 2 ?

Nnesha (nnesha):

 2 is a base number at front of log became exponent of k